I was asked to solve for y.... x+4y=4.
the answer I was given is y= (x/4)-1. How did they get this? Step by step please... I'm new to this! Thank you!

Answer by Alan3354(69443)

x+4y=4
Subtract x
4y = 4 - x
Divide by 4
y = 1 - (x/4)
-------------
The answer you were given is wrong.